OpenStudy (anonymous):

How do I find the slope of 12x + 4y = 24 ?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

4y = 12x + 24

OpenStudy (anonymous):

4/4y = 12/4 + 24/4

OpenStudy (anonymous):

y=3x + 6

OpenStudy (anonymous):

had trouble with these kinds of problems to so thats all the steps I know

OpenStudy (anonymous):

Thankss .

OpenStudy (anonymous):

No problem, I had these kinds of homework problems last night and I didn't know the rest of the steps, just how to simplify it.

OpenStudy (anonymous):

I got confused ! &- I still am .

OpenStudy (anonymous):

Ask your teacher and maybe he could help you. Afterall it is what they're there for.

OpenStudy (anonymous):

A line has a slope of m = 12 , and the point ( -2, 8 ) lies on the line . What is the equation of the line ?

OpenStudy (anonymous):

you can also use the fact that the slope of \[ax+by=c\] is \[-\frac{a}{b}\]